### v1

@@ Abstract.
For a positive integer $k$, the $k$-cut complex of a graph $G$ is the simplicial complex whose facets are the $(|V(G)|-k)$-subsets $\sigma$ of the vertex set $V(G)$ of $G$ such that the induced subgraph of $G$ on $V(G)\setminus\sigma$ is disconnected. These complexes first appeared in the master thesis of Denker and were further studied by Bayer et al. in [Topology of cut complexes of graphs, SIAM Journal on Discrete Mathematics, 2024]. In the same article, Bayer et al. conjectured that for $k\geq 3$, the $k$-cut complexes of squared cycle graphs are shellable. Moreover, they also conjectured about the Betti numbers of these complexes when $k=3$. In this article, we prove these conjectures for $k=3$.
@@ 1. Introduction
Inspired from Fröberg’s theorem (Theorem 1.1) and (1.1), Denker [5] introduced $k$-cut complexes of graphs which are a generalization of $\Delta_{2}(G)=\mathsf{Cl}(G)^{\vee}$. In [2], Bayer et al. introduced an another generalization of $\Delta_{2}(G)$, the total $k$-cut complexes. For $k\geq 1$, the total $k$-cut complex of a graph $G$, denoted as $\Delta_{k}^{t}(G)$, is the simplicial complex whose facets are $\sigma\subseteq V(G)$ such that $|\sigma|=|V(G)|-k$ and the induced subgraph $G[V(G)\setminus\sigma]$ does not contain any edge. In particular, $\Delta_{2}(G)=\Delta_{2}^{t}(G)$ for any graph $G$.
For $n\geq 3$, the squared cycle graph $W_{n}$ is a graph where the set of vertices $V(W_{n})=\{0,1,2,\ldots,n-1\}$ and the set of edges $E(W_{n})=\{\{i,i+1\ (\text{mod}\ n)\},\{i,i+2\ (\text{mod}\ n)\}:0\leq i\leq n-1\}$. For $n\geq 5$, squared cycle graphs $W_{n}$ are also $4$-regular circulant graphs and hence Cayley graphs of the cyclic group on $n$ elements (for the definition of circulant graphs see Section 2). In [2, Proposition 4.19], authors studied the total $2$-cut complexes of squared cycle graphs and proved that these complexes are homotopy equivalent to wedge of spheres of dimension lower than the dimension of $\Delta_{2}^{t}(W_{n})=\Delta_{2}(W_{n})$. Therefore, the complexes $\Delta_{2}(W_{n})$ are not shellable. Further, based on their observations supported by Sage calculations (for $n\leq 13$ and $k\leq 5$), they conjectured that $\Delta_{k}(W_{n})$ are shellable for $k\geq 3$.
@@
For $k\geq 3$, the cut complex $\Delta_{k}(W_{n})$ is shellable for $n\geq k+6$. For $k=3$ and $n\geq 9,$ the Betti numbers are $\binom{n-4}{2}-9=\{1,6,12,19,27,\dots\}$.
@@
Let $n\geq 9$. Then the $3$-cut complex $\Delta_{3}(W_{n})$ of a squared cycle graph is shellable. Moreover,
$\Delta_{3}(W_{n})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$
@@
This paper is organized as follows: In Section 2, we provide the necessary preliminaries related to graph theory and simplicial complexes. In Section 3, we give the proof of Theorem 1.4. This section is divided into two subsections. To prove Theorem 1.4, we define an order in the facets of $\Delta_{3}(W_{n})$ and in Section 3.1, we show that this order is a shelling order. In Section 3.2, we characterize and count the number of spanning facets for the shelling order to conclude the number of spheres appearing in the wedge in the homotopy type of $\Delta_{3}(W_{n})$.
@@ 2.1. Graph
A graph $G$ is a pair $(V(G),E(G))$, where $V(G)$ is the set of vertices of $G$ and $E(G)\subset\binom{V(G)}{2}$ denotes the set of edges. If $\{x,y\}\in E(G)$, it is also denoted by $x\sim y$ and we say that $x$ is adjacent to $y$. A subgraph $H$ of $G$ is a graph with $V(H)\subset V(G)$ and $E(H)\subset E(G)$. For a subset $U\subseteq V(G)$, the induced subgraph $G[U]$ is the subgraph whose set of vertices is $V(G[U])=U$ and the set of edges is $E(G[U])=\{\{a,b\}\in E(G)\ |\ a,b\in U\}$.
@@
Let $n\geq 2$ be a positive integer and $S\subset\{1,2,\ldots,n-1\}$. The circulant graph $C_{n}(S)$ is the graph whose set of vertices is $V(C_{n}(S))=\{0,1,2,\ldots,n-1\}$ and any two vertices $x$ and $y$ are adjacent if and only if $x-y\ (\text{mod \ $n$})\in S\cup-S$, where $-S=\{n-a\ |\ a\in S\}$. Circulant graphs are also Cayley graphs of $\mathbb{Z}_{n}$, the cyclic group on $n$ elements. For $n\geq 3,$ the squared cycle graph $W_{n}$ on $n$ vertices, is a circulant graph $C_{n}(S),$ where $S=\{1,2\}.$
@@ 2.2. Simplicial complex
A finite abstract simplicial complex $\Delta$ is a collection of finite sets such that if $\tau\in\Delta$ and $\sigma\subset\tau$, then $\sigma\in\Delta$. The elements of $\Delta$ are called simplices of $\Delta$. The dimension of a simplex $\sigma$ is equal to $|\sigma|-1$. The dimension of an abstract simplicial complex is the maximum of the dimensions of its simplices. The 00-dimensional simplices are called vertices of $\Delta$, and the set of vertices of $\Delta$ is denoted by $V(\Delta)$. If $\sigma\subset\tau$, we say that $\sigma$ is a face of $\tau$. If a simplex has dimension $d$, it is said to be $d$-dimensional. A simplex that is not a face of any other simplex is called a maximal simplex or facet. The set of maximal simplices of $\Delta$ is denoted by $M(\Delta)$. A simplicial complex is called pure $d$-dimensional, if all of its maximal simplices are of dimension $d$.
In this article, we consider any simplicial complex as a topological space, namely, its geometric realization. For the definition of geometric realization, we refer to the book [10] by Kozlov. For terminologies of algebraic topology used in this article, see [8].
@@ 2.3. Shellability
[…]
@@
Definition 2.5 ([10, Section 12.1]).
Let $\Delta$ be a shellable simplicial complex, and let $F_{1},F_{2},\ldots,F_{t}$ be a shelling on $\Delta$. Then $F_{j}$ $(1<j\leq t)$ is said to be a spanning facet for the given shelling order if for each $\lambda\in F_{j}$, there exists $r<j$ such that $F_{r}\cap F_{j}=F_{j}\setminus\{\lambda\}$.
@@
In this section, we prove the main result (Theorem 1.4) of this article. Throughout the section, we fix $n\geq 9$. Our aim is to define an order $\prec$ in the facets of $\Delta_{3}(W_{n})$, which provides a shelling order.
@@
We arrange the elements of $V(W_{n})=\{0,1,2,\ldots,n-1\}$ into an ordered set $\mathcal{O}:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$, where $\alpha_{t}=m+(-1)^{t-1}\lfloor(t/2)\rfloor\ (\text{mod}\ n)$ $\forall$ $1\leq t\leq n$ and $\lfloor(t/2)\rfloor$ denotes the greatest integer less than or equal to $t/2$. Hence
@@
We define an order $<_{\mathcal{O}}$ in the elements of $V(W_{n})$ as follows: For any $x,y\in V(W_{n})$, we say that $x<_{\mathcal{O}}y$ if and only if $x=\alpha_{s}$ and $y=\alpha_{t}$ for some $1\leq s<t\leq n$.
@@
Throughout this section, we denote the complement of a set $X\subseteq V(W_{n})$ by $X^{c}$.
Let $\mathcal{T}$ be the set of all $(n-3)$-subsets of $V(W_{n})$. We partition $\mathcal{T}$ into disjoint sets $\mathcal{T}_{s}$ as follows: Let $\mathcal{T}_{1}:=\{\tau\in\mathcal{T}:\alpha_{1}\in\tau^{c}\}.$ For $s>1$, define
@@
Let $M(\Delta_{3}(W_{n}))$ be the set of facets of $\Delta_{3}(W_{n})$. For any $F\in M(\Delta_{3}(W_{n}))$, we have $|F|=n-3$ and $W_{n}[F^{c}]$ is disconnected. Clearly, $M(\Delta_{3}(W_{n}))\subset\mathcal{T}=\bigsqcup\mathcal{T}_{s}$.
@@
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Observe that since $W_{n}[F^{c}]$ is disconnected and $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$, we get $\alpha_{s}\notin\{0,1,n-1\}$. Hence $22\leq\alpha_{s}\leq n-$ 2.
For each $s$, we define an order $<_{\mathcal{T}_{s}}$ in the elements of $\mathcal{T}_{s}$ as follows: if $F$, $F^{\prime}$in $\mathcal{T}_{s}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s^{\prime}_{1},s^{\prime}_{2}\}$, then $F<_{\mathcal{T}_{s}}F^{\prime}$ if and only if either $s_{1}<s^{\prime}_{1}$ or, $s_{1}=s^{\prime}_{1}$ and $s_{2}<s^{\prime}_{2}$.
Now using the order $<_{\mathcal{T}_{s}}$, we define an order $\ll$ in the elements of $M(\Delta_{3}(W_{n}))$ as follows: Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$. Then $F\ll F^{\prime}$ if and only if either $s<t$ or, $s=t$ and $F<_{\mathcal{T}_{s}}F^{\prime}$.
To get our desired shelling order $\prec$, we slightly modify the order $\ll$ by “displacing” some elements of the poset $(M(\Delta_{3}(W_{n})),\ll)$ and get a new poset $(M(\Delta_{3}(W_{n})),$ $\prec$)≺ ). For this, we first define a subset $\mathcal{D}$ of $M(\Delta_{3}(W_{n}))$ as follows: A facet $F\in\mathcal{D}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies any of the following conditions:

### v2

@@ Abstract.
For a positive integer $k$, the $k$-cut complex of a graph $G$ is the simplicial complex whose facets are the $(|V(G)|-k)$-subsets $\sigma$ of the vertex set $V(G)$ of $G$ such that the induced subgraph of $G$ on $V(G)\setminus\sigma$ is disconnected. These complexes first appeared in the master thesis of Denker and were further studied by Bayer et al. in [Topology of cut complexes of graphs, SIAM Journal on Discrete Mathematics, 38(2):1630-1675, 2024]. In the same article, Bayer et al. conjectured that for $k\geq 3$, the $k$-cut complexes of squared cycle graphs are shellable. Moreover, they also conjectured about the Betti numbers of these complexes when $k=3$. In this article, we prove these conjectures for $k=3$.
Key words and phrases:
Cut complex, Shellability, Squared cycle graphs, Homotopy
@@ 1. Introduction
Inspired from Fröberg’s theorem (Theorem 1.1) and (1.1), Denker [5] introduced $k$-cut complexes of graphs which are a generalization of $\Delta_{2}(G)=\mathsf{Cl}(G)^{\vee}$. In [2], Bayer et al. introduced another generalization of $\Delta_{2}(G)$, the total $k$-cut complexes. For $k\geq 1$, the total $k$-cut complex of a graph $G$, denoted as $\Delta_{k}^{t}(G)$, is the simplicial complex whose facets are $\sigma\subseteq V(G)$ such that $|\sigma|=|V(G)|-k$ and the induced subgraph $G[V(G)\setminus\sigma]$ does not contain any edge. In particular, $\Delta_{2}(G)=\Delta_{2}^{t}(G)$ for any graph $G$.
For a graph $G$ and a positive integer $p$, the $p$-th power graph of $G$ is a graph $G^{p}$ where the set of vertices $V(G^{p})=V(G)$ and any $\{u,v\}$ is an edge in $G^{p}$ if and only if there exists a path between $u$ and $v$ of length at most $p$ in $G$ (here, the length of a path is the number of edges in the path). Clearly $G^{1}=G$. For $n\geq 3$, the cycle graph $C_{n}$ on $n$ vertices is a graph where the set of vertices $V(C_{n})=\{0,1,2,\ldots,n-1\}$ and the set of edges $E(C_{n})=\{\{i,i+1\ (\text{mod}\ n)\}:0\leq i\leq n-1\}$. Observe that the $p$-th powered cycle graph $C_{n}^{p}$ on $n$ vertices is a graph with $V(C_{n}^{p})=V(C_{n})$ and the set of edges $E(C_{n}^{p})=\{\{i,i+j\ (\text{mod}\ n)\}:0\leq i\leq n-1\text{ and }1\leq j\leq p$}. For $n\geq 2p+1$, powered cycle graphs $C_{n}^{p}$ are also $2p$-regular circulant graphs and hence Cayley graphs of the cyclic group on $n$ elements (for the definition of circulant graphs see Section 2).
It can be easily observed that for a graph $G$, $\Delta_{1}(G)$ is void. In [2], authors studied the total $2$-cut complexes of the cycle graphs $C_{n}$ [2, Theorem 4.15] and the squared cycle graphs $C_{n}^{2}$ [2, Proposition 4.19]. They proved that these complexes (if nonvoid) are homotopy equivalent to wedges of spheres of dimension lower than the dimension of $\Delta_{2}^{t}(C_{n}^{p})=\Delta_{2}(C_{n}^{p})$ for $p=1,2$. Therefore, the complexes $\Delta_{2}(C_{n})$ and $\Delta_{2}(C_{n}^{2})$ are not shellable. In [1, Proposition 7.11], authors proved that for $k\geq 3$, $\Delta_{k}(C_{n})$ is shellable for $n\geq k+1$. Further, based on their observations supported by Sage calculations (for $n\leq 13$ and $k\leq 5$), they conjectured that for $k\geq 3$, $\Delta_{k}(C_{n}^{2})$ are shellable for $n\geq k+6$.
@@
For $k\geq 3$, the cut complex $\Delta_{k}(C_{n}^{2})$ is shellable for $n\geq k+6$. For $k=3$ and $n\geq 9,$ the Betti numbers are $\binom{n-4}{2}-9=\{1,6,12,19,27,\dots\}$.
@@
Let $n\geq 9$. Then the $3$-cut complex $\Delta_{3}(C_{n}^{2})$ of a squared cycle graph is shellable. Moreover,
$\Delta_{3}(C_{n}^{2})\simeq\bigvee\limits_{\binom{n-4}{2}-9}\mathbb{S}^{n-4}.$
@@
We continued our investigation to analyze the $k$-cut complex $\Delta_{k}(C_{n}^{p})$ for powered cycle graphs $C_{n}^{p}$ (for small values of $k$, $p$ and $n$) and computed their homology groups (with coefficient $\mathbb{Z}$) using SageMath (see Tables 1-5 in Section 4). By observing Table 1 and the fact that the complexes $\Delta_{2}(C_{n})$ and $\Delta_{2}(C_{n}^{2})$ (if nonvoid) are not shellable, we conjecture that if $\Delta_{2}(C_{n}^{p})$ is nonvoid, then it is not shellable for $p\geq 3$ (4.1).
It is proved in [1, Proposition 7.11] that $\Delta_{3}(C_{n})$ is shellable for $n\geq 5=4+1$ (for $n\leq 4$, $\Delta_{3}(C_{n})$ is void). In Theorem 1.4, we have proved that $\Delta_{3}(C_{n}^{2})$ is shellable for $n\geq 9=4.2+1$. Therefore, for $p\in\{1,2\}$, $\Delta_{3}(C_{n}^{p})$ is shellable for $n\geq 4p+1$. Based on these results and Table 2, we conjecture that $\Delta_{3}(C_{n}^{p})$ is shellable for all $p\geq 3$ and $n\geq 4p+1$ (4.2).
Tables 3, 4 and 5 bring up the question of whether the complexes $\Delta_{k}(C_{n}^{p})$ (if nonvoid) are not shellable for $k\geq 4$ and $p\geq 3$. We also make a conjecture about the homotopy types of the $k$-cut complexes $\Delta_{k}(C_{n}^{2})$ of squared cycles (4.3).
This paper is organized as follows: In Section 2, we provide the necessary preliminaries related to graph theory and simplicial complexes. In Section 3, we give the proof of Theorem 1.4. This section is divided into two subsections. To prove Theorem 1.4, we define an order in the facets of $\Delta_{3}(C_{n}^{2})$ and in Section 3.1, we show that this order is a shelling order. In Section 3.2, we characterize and count the number of spanning facets for the shelling order to conclude the number of spheres appearing in the wedge in the homotopy type of $\Delta_{3}(C_{n}^{2})$. Finally, in Section 4, based on our SageMath data, we propose several conjectures and questions about the shellability and homotopy types of the complexes $\Delta_{k}(C_{n}^{p})$.
@@ 2.1. Graph
A graph $G$ is a pair $(V(G),E(G))$, where $V(G)$ is the set of vertices of $G$ and $E(G)\subseteq\binom{V(G)}{2}$ denotes the set of edges. If $\{x,y\}\in E(G)$, it is also denoted by $x\sim y$ and we say that $x$ is adjacent to $y$. A subgraph $H$ of $G$ is a graph with $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. For a subset $U\subseteq V(G)$, the induced subgraph $G[U]$ is the subgraph whose set of vertices is $V(G[U])=U$ and the set of edges is $E(G[U])=\{\{a,b\}\in E(G)\ |\ a,b\in U\}$.
For $x,y\in V(G)$, a path from $x$ to $y$ is a sequence $xv_{0}\ldots v_{n}y$ of vertices of $G$ such that $x\sim v_{0},v_{n}\sim y$ and $v_{i}\sim v_{i+1}$ for all $10\leq i\leq n-$ 1. The number of edges in a path is the length of the path.
@@
Let $n\geq 2$ be a positive integer and $S\subset\{1,2,\ldots,n-1\}$. The circulant graph $C_{n}(S)$ is the graph whose set of vertices is $V(C_{n}(S))=\{0,1,2,\ldots,n-1\}$ and any two vertices $x$ and $y$ are adjacent if and only if $x-y\ (\text{mod \ $n$})\in S\cup-S$, where $-S=\{n-a\ |\ a\in S\}$. Circulant graphs are also Cayley graphs of $\mathbb{Z}_{n}$, the cyclic group on $n$ elements.
@@ 2.2. Simplicial complex
A finite abstract simplicial complex $\Delta$ is a collection of finite sets such that if $\tau\in\Delta$ and $\sigma\subset\tau$, then $\sigma\in\Delta$. The elements of $\Delta$ are called simplices of $\Delta$. If $\sigma\subset\tau$, we say that $\sigma$ is a face of $\tau$. The dimension of a simplex $\sigma$ is equal to $|\sigma|-1$. The dimension of an abstract simplicial complex is the maximum of the dimensions of its simplices. The 00-dimensional simplices are called vertices of $\Delta$, and the set of vertices of $\Delta$ is denoted by $V(\Delta)$. If a simplex has dimension $d$, it is said to be $d$-dimensional. An abstract simplicial complex which is an empty collection of sets is called the void abstract simplicial complex, and is denoted by $\emptyset$.
The boundary of a $d$-dimensional simplex $\sigma$ is the simplicial complex, consisting of all faces of $\sigma$ of dimension $\leq d-1$ and it is denoted by $Bd(\sigma).$ A simplex that is not a face of any other simplex is called a maximal simplex or facet. The set of maximal simplices of $\Delta$ is denoted by $M(\Delta)$. A simplicial complex is called pure $d$-dimensional, if all of its maximal simplices are of dimension $d$. A subcomplex $\Delta^{\prime}$ of $\Delta$ is a simplicial complex such that $\sigma\in\Delta^{\prime}$ implies $\sigma\in\Delta$.
In this article, we consider any simplicial complex as a topological space, namely, its geometric realization. For the definition of geometric realization, we refer to the book [10] by Kozlov. For terminology of algebraic topology used in this article, see [8].
@@ 2.3. Shellability
[10, Section 12.1] A simplicial complex $\Delta$ is called shellable if its facets can be arranged in a linear order $F_{1},F_{2},\ldots,F_{t}$ in such a way that for all $2\leq j\leq t$, the subcomplex $(\bigcup_{i=1}^{j-1}F_{i})\cap F_{j}$ is pure and of dimension $|F_{j}|-2$.
@@
A facet $F_{k}$ $(1<k\leq t)$ is called spanning with respect to the given shelling order if $Bd(F_{k})\subseteq\bigcup\limits_{i=1}^{k-1}F_{i}$. It is easy to check that $F_{k}$ is a spanning facet if for each $\lambda\in F_{k}$, there exists $r<k$ such that $F_{r}\cap F_{k}=F_{k}\setminus\{\lambda\}$.
@@
In this section, we prove the main result (Theorem 1.4) of this article. Throughout the section, we fix $n\geq 9$. Our aim is to define an order $\prec$ in the facets of $\Delta_{3}(C_{n}^{2})$, which provides a shelling order.
@@
We arrange the elements of $V(C_{n}^{2})=\{0,1,2,\ldots,n-1\}$ into an ordered set $\mathcal{O}:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$, where $\alpha_{t}=m+(-1)^{t-1}\lfloor(t/2)\rfloor\ (\text{mod}\ n)$ $\forall$ $1\leq t\leq n$ and $\lfloor(t/2)\rfloor$ denotes the greatest integer less than or equal to $t/2$. Hence
@@
We define an order $<_{\mathcal{O}}$ in the elements of $V(C_{n}^{2})$ as follows: For any $x,y\in V(C_{n}^{2})$, we say that $x<_{\mathcal{O}}y$ if and only if $x=\alpha_{s}$ and $y=\alpha_{t}$ for some $1\leq s<t\leq n$.
@@
Throughout this section, we denote the complement of a set $X\subseteq V(C_{n}^{2})$ by $X^{c}$.
Let $\mathcal{T}$ be the set of all $(n-3)$-subsets of $V(C_{n}^{2})$. We partition $\mathcal{T}$ into disjoint sets $\mathcal{T}_{s}$ as follows: Let $\mathcal{T}_{1}:=\{\tau\in\mathcal{T}:\alpha_{1}\in\tau^{c}\}.$ For $s>1$, define
@@
Let $M(\Delta_{3}(C_{n}^{2}))$ be the set of facets of $\Delta_{3}(C_{n}^{2})$. For any $F\in M(\Delta_{3}(C_{n}^{2}))$, we have $|F|=n-3$ and $C_{n}^{2}[F^{c}]$ is disconnected. Clearly, $M(\Delta_{3}(C_{n}^{2}))\subset\mathcal{T}=\bigsqcup\mathcal{T}_{s}$.
@@
Let $F\in M(\Delta_{3}(C_{n}^{2}))$ such that $F\in\mathcal{T}_{s}$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Observe that since $C_{n}^{2}[F^{c}]$ is disconnected and $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$, we get $\alpha_{s}\notin\{0,1,n-1\}$. Hence $22\leq\alpha_{s}\leq n-$ 2.
Definition 3.4.
Define an order $\ll$ in the elements of $M(\Delta_{3}(C_{n}^{2}))$ as follows: Let $F,F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ such that $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and $F^{\prime c}=\{\alpha_{t}\}\sqcup\{t_{1},t_{2}\}$. Then $F\ll F^{\prime}$ if and only if any one of the following conditions is true:
  1. (i)
    If $s=t$ i.e., $\alpha_{s}=\alpha_{t}$, then either $s_{1}<t_{1}$ or, $s_{1}=t_{1}$ and $s_{2}<t_{2}$.
  2. (ii)
    If $s\neq t$ i.e., $\alpha_{s}\neq\alpha_{t}$, then $s<t$ i.e., $\alpha_{s}<_{\mathcal{O}}\alpha_{t}$.
It is immediate that Definition 3.4 defines a total order on $M(\Delta_{3}(C_{n}^{2}))$.
To get our desired shelling order $\prec$, we slightly modify the order $\ll$ by “displacing” some elements of the poset $(M(\Delta_{3}(C_{n}^{2})),\ll)$ and get a new poset $(M(\Delta_{3}(C_{n}^{2})),$ $\prec$)≺ ). For this, we first define a subset $\mathcal{D}$ of $M(\Delta_{3}(C_{n}^{2}))$ as follows: A facet $F\in\mathcal{D}$ if and only if there exists $s$ such that $F\in\mathcal{T}_{s}$ and $F$ satisfies any of the following conditions: